OpenStudy (17bohtal):

Solve the equation 9+n/5=19

OpenStudy (anonymous):

n = 50

OpenStudy (anonymous):

9 + 50/5 = 19 9 + 10 = 19

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

i can explain it in a different way if you want me to

OpenStudy (anonymous):

9+n/5=19 -9 -9 n/5=10 *5 *5 n=50

OpenStudy (17bohtal):

Thank you!

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

yout welcome

OpenStudy (anonymous):

@joshhfers explained it better though